Grade 12th PassMechanics

A piston of mass M = 3kg and radius R = 4cm has a hole into which a thin pipe of radius r = 1cm is inserted. The piston can enter a cylinder tightly and without friction, and initially it is at the bottom of the cylinder. 750gm of water is now poured into the pipe so that the piston & pipe are lifted up as shown. Find the height H of water in the cylinder and height h of water in the pipe.

Koolman
10 Years agoGrade 12th Pass

1 Answer

Garvita Singhvi
10 Years ago
750gm=750*106 m3 =M   ….........................(1gm=1cm3=106m3)
3g=h∫g(42-12)π                           …...….........(∫=dencity)
M=π(42H+12h)
solving
H=11/(32π)
h=2/π
 
I hope it helps